OpenStudy (smile.89745):

Can you please help me!!! Consider the following system of equations: y = −x + 2 y = 3x + 1 Which description best describes the solution to the system of equations? Line y = −x + 2 intersects line y = 3x + 1. Lines y = −x + 2 and y = 3x + 1 intersect the x-axis. Lines y = −x + 2 and y = 3x + 1 intersect the y-axis. Line y = −x + 2 intersects the origin.

OpenStudy (smile.89745):

@sammixboo @SkyVoltage @Yuii @drawwithapurpose @Anna1448

OpenStudy (smile.89745):

I have no idea about this and i need help

OpenStudy (skyvoltage):

hmmmm give me a minute ok

OpenStudy (smile.89745):

Okay i have 2 more after

OpenStudy (drawwithapurpose):

i believe it's a

OpenStudy (skyvoltage):

ok so its not b or d

OpenStudy (skyvoltage):

and it wont cross on y axis so yes A is correct

OpenStudy (smile.89745):

Okay 2 more

OpenStudy (skyvoltage):

ok

OpenStudy (smile.89745):

Equation R: −3y = −3x – 9 Equation S: y = x + 3 Which of the following best describes the solution to the given set of equations? No solution One solution Infinite solutions Two solutions

OpenStudy (smile.89745):

and this one Consider the following pair of equations: y = x + 4 y = −2x − 2 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

OpenStudy (skyvoltage):

D

OpenStudy (smile.89745):

what about the last one and thank you. You know that first one for sure right?

OpenStudy (skyvoltage):

yea

OpenStudy (smile.89745):

okay what about the last one

OpenStudy (skyvoltage):

@drawwithapurpose hey can you help me out

OpenStudy (skyvoltage):

im not good at explaining

OpenStudy (drawwithapurpose):

yeah sure just a sec

OpenStudy (drawwithapurpose):

y = x; 4y=−2x−2 Solve y = x for y: Substitute x for y in 4y=−2x−2: 4y=−2x−2 4x=−2x−2 4x+2x=−2x−2+2x(Add 2x to both sides) 6x=−2 6x/6= −2/6 (Divide both sides by 6) x=−1/3 Substitute for x in y = x: y = −1/3 there ya go x3

OpenStudy (smile.89745):

Okay thanks and your sure?

OpenStudy (drawwithapurpose):

positive :3

OpenStudy (smile.89745):

Thank you to the both of you so much

OpenStudy (drawwithapurpose):

no problem

OpenStudy (skyvoltage):

welcome

OpenStudy (smile.89745):

Couuld this be right? 1.)Replace for y --->>> (y=x+4) --->>> x+4=-2x-2 2.) x + 2x + 4 = -2x + 2x - 2 3.) 3x + 4 = -2 4.) 3x + 4 -4 = -2 - 4 5.) 3x = -6 6.) 3x/3 = -6/3 7.) x = 2 Solving for y 1.) y = x + 4 2.) y = 2 + 4 3.) 4 + 2 = 6 4.) y = 6 Solution would be (-2,6)

OpenStudy (smile.89745):

@drawwithapurpose

OpenStudy (drawwithapurpose):

mhm that would be another way to do it but you would need to make -6/3 -2 instead of 2 because when you multiply a positive and negative its always negative

OpenStudy (smile.89745):

Okay thanks again